Startup Night 7 - 20170908 UT

Observers: JHill, BRothberg, A.Cardwelll (Tucson)
Telescope Operator: G.Bechetti (LBTO)
AO Support: D.Miller
Telescope Support: KNewton (LBTO)
Software Support: CBiddick, KSummers (Tucson)
Instrument Support: J.Morris (LBTO)
Engineering: TMcMahon


Plan


Log

UT 01:33 - Started LBCs

UT 01:35 - GEIRS1 Error on LUCI-1 updated IT 6762, restarted GEIRS1 service

UT 01:52 - Taking 10 biases LBC each side

AGW3 PEPSIPOL Part 1 - Initial Configuration on-sky

Steps:
  Pointing - done with old pointing model
  Field Orientation - 90% done
  Acquisition off-axis
  Off-Axis Guiding
  wfs
  Probe Centering on rotator axis

UT 02:00 - Updated AzCam port numbers for AGW3.cfg and AGW4.cfg. After that, GSCL is able to initialize AzCam OK.

UT 02:02 - ACQUIRE preset hits M5 star in the corner (using old pointing model from pre-2017) or was it a ghost? IE=-111 CA=-80

Eyeball Coma correction -2300 Z7 Focus -0.9 mm

UT 02:15 ACQUIRE preset on pointing star (HD 195557)

UT 02:20 TRACK preset to alignment field from Ilya - ADS 9371 AB.

UT 02:35 M5 star near alignment field Focus -3.7 mm for stage =31.0 IE -122 CA=-73 (why did focus change?)

UT 02:xx TRACK preset to alignment field from Ilya - ADS 9371 AB. Secondary star was lower right on guide camera at commanded PA of 0.0.

IE=-136 CA=-73

03:20 Change from LEFTIAA=0 to LEFTIAA=+90 and restart PCS

IE=-136 CA=-83

03:52 ACQUIRE preset to BD+30 3779C IE=-143 CA=-116 (secondary star was to the bottom on guide camera - but maybe that was for parallactic mode)

Why is focus (AGW-Z) going to 0.0 sometimes? Because it is a TRACK preset and the probe is parked.

UT 04:11 ACQUIRE IE=-147 CA=-90 (secondary star is to the left in parallactic preset)

after centering on the guide probe - it needs +10 IE to get star onto the pinhole

UT 04:15 Change LEFTIAA from +90 to +180 deg.

UT 04:24 Change LEFTIAA from +180 to +135

UT 04:27 Change LEFTIAA from +135 to +225 - here the RADEC offsets move in the direction that Ilya expects. or perhaps 180 opposite. Rotator angle is at 120 deg.

LBTI AO Wake-up

UT 04:35 Switchover to LBTI for AO and M3 alignment with Amali and Phil. LBTI general pointing was confirmed, but other work was postponed until later.

AGW3 PEPSIPOL Part2

UT 05:30 Switch back to PEPSIPOL for more AGw alignment fun.

UT 05:42 ACQUIRE preset to BD+30 3779C

IP of pepsib is 192.168.164.19 for vnc connection

IE=-144 CA=-80 (centered on the secondary star)

Change IAA from 225 to 45 deg. Rotator angle is 303 at PosAng =0

IE=-145 CA=-74 (centered between the two stars)

UT 06:15 cannot move probe to SFP -23,-55 mm - ACQUIRE failed for a guide star on the east side.

UT 06:20 changing GCS SFP_rotation from 179 to 90 degrees changed nothing (probe still wants -23,-55)

06:23 moved the +45 from the variable LEFTIAA to LEFTZEROPOINT

This changed the cannot move probe error to SFP +22,-55 when ACQUIRE failed. This confirms that the two variables are NOT equivalent as we thought last week.

06:29 changed LEFTXYFLIP from false/false to false/true - this made the rotator turn 180, why? Michele's explanation is that we have changed the instrument transformation in the kernel.

06:32 cannot move probe to SFP -23,-55 ACQUIRE failed

06:35 moved the +45 from ZEROPOINT back to IAA (again the rotator turns but by 90 !!!)

cannot move probe to SFP -55,-23 ACQUIRE failed

06:40 changing stars to one approaching the meridian 27 PSC - THIS ACQUIRE preset did not fail for an off-axis guide star (-92,+98 which is a valid +Y value). IE=-136 CA=-77

06:50 changed IAA from +45 to +135.

06:52 ACQUIRE succeeded at SFP +98,+92 but we didn't see the off-axis star on the guider camera.

06:54 changed to 54 Acq on merdian. ACQUIRE did not fail at SFP -14, 115, but did not find the off-axis star. IE -145,CA=-75 Visual inspection says the instrument/rotator orientation is correct, and the RADEC offsets seem to move the star the right way.

This is where we stopped without seeing the star on the off-axis guide star. Either we are off by a small amount in angle, or we have an X-flip problem with the guide probe.

INSTRUMENT      PEPSIPOL
FOCALSTATION      directGregorian
LEAKRATE      0.5
FLOOR         0.4
POWER         1
RANGE         0.25
LEFTPOINTINGMODEL   SX_Gregorian.20170414.ptmod
LEFTFOCALLENGTH         123421.4
LEFTPOINTINGORIGIN   0.0 0.0
LEFTXYFLIP      false true
LEFTIAA         135.0
LEFTPAI         0.0
LEFTWAVELENGTH      0.500
LEFTZEROPOINT      0.0
# Rotation Scale negative for Direct focus rotators
LEFTSCALE      -1.0
LEFTGUIDEWAVELENGTH     0.630
LEFTAOREFWAVELENGTH     0.750

UT 07:02 reconfigure to LUCI/LUCI

Testing Dave T. Pointing Model from 20170907 UT

UT 07:15 - will move to various stars to check model

UT 07:22 - pointing preset

UT 07:26 - PCS crashed - it doesn't read the binocular file format.

UT 07:38 - PCS fixed, reformatted file. Moving to various pointing stars at various elevations.

UT 07:39 - moving to M5 pointing star near Stone 0 "and away we go! - wabba lubba dub dub!"

UT 07:44 - Dave T's new pointing model is AWESOME!

AGW 1 & 2 Transform Test (Stone O field)

-will use Stone_O_0.list, Stone_O_p10.list, Stone_O_n10.list

UT 07:52 - IDL problem, some syntax error crashes transform_collect - issue appears to be related to a recent library update needed to analyze field aberration data. Effect is on lbto account on obs N workstations.

UT 08:02 - atempting to fix by copying older library back over to directory

UT 08:08 - co pointing failed on left side, right side hung, Doug M. sending a binocular preset GS0

UT 08:19 - did a binocular preset to GS=0 to avoid breaking copointing limit -starting collection both sides

-make sure to exit IDL before moving to Stone_O_p10 and Stone_O_n10

UT 08:25 - Barry forgot to configure LUCI 1 & 2 for N375 with N30 field stop & BrGamma - 5sec DIT, exited IDL so Doug's update to movement to 20" will kick in to effect

BE SURE TO TAKE A 5SEC BLIND,BLIND for L1 and L2

UT 08:32 - Restarted Stone_O_0.list L1# = 0039 L2# = 0009

UT 08:35 - GEIRS2 Sync error couldn't take image -stopping

UT 08:37 - presets cancelled, only GS11 (1st in list was done, so restarting whole thing), L1#=0041-0053, L2#=0011-0023

UT 08:58 - SX finished

UT 09:00 - DX finished

UT 09:01 - Running Stone_O_p10.list L1#=0054 L2#=0024

UT 09:04 - GEIRS L1 aborted - started early enough that we hadn't collected a lot of data - restarting GEIRS1 service ignore

UT 09:06 - took test exposure after restart and now GEIRS1 stuck in saving - stuck on 54 and 54. Now restarting LUCI-1 software

UT 09:13 - turning off LUCI-1 only and running cleanup also turned off LUCI-2, restarted all LUCI softwware - now both GEIRS have error

UT 09:15 - also initialized all - took manual exposures through GEIRS windows those were ok - error because there are two #54s on LUCI-1, LUCI-2 took exposure with READOUT MANAGER OK.

Calling Tom Sargent to fix duplicate file - fixed by logging in as readou1 and rm'ing file

UT 09:27 - Resting shells at elevation 30 - moving to next set in field - called Doug M. who said should be able to take transform data at this elevation with shells rested

UT 09:31 - Restarted at Stone_O_p10.list L1# = 0055 L2#=0027

UT 09:36 - this isn't working, can't collimate enough to get anything round - moving back up to zenith to restore shells

LUCI-Bino G210 N375 Test

UT 09:47 - will point and collimate before moving to NGC 828 for test

UT 10:12 - some issues with AdSecs collimating, finally OK, now sending preset

UT 10:17 - error on instrument config luci1 - not clear what the actual problem is, resent config, now skipping to next line

UT 10:21 - FM4 is ~180 pixels below (Y~1870) - this is an issue because i'm testing a 60" dither

UT 10:23 - will do a manual alignment with AFC tab once 3 acquisition images are done

UT 10:28 - i cancelled presets and resending acquisition with fixed FM4, reason is there is rotation in slit projection and with 0.5" slit alignment might be affected by moving FM before slit alignment

UT 10:40 - started spectroscopy NGC 828

UT 11:00 - finsihed target, moving to telluric HD 12381

UT 11:13 - finished telluric

Target test for Barry

UT 11:14 - moving to G200 spectra zJspec of FF0050-0039 (mid z ULIRG - little time left will try to get some data)

UT 11:16 - need to check pointing on SX

UT 11:20 - resending preset

UT 11:21 - SX preset failed again.

UT 11:25 - moved to FF0139+0115

UT 11:29 - LUCI-2 MOS error while putting mask from turnout to FPU, Rotator = 328, elevation 50 degrees

14010962 2017-09-08 11:28:39.0770 ERROR *** Luci TWO TWO_MOSUnit

Error while releasing mask! Moved -5658 steps without reaching limit switch where -8000 steps should be enough.
UT 11:32 - Called Dave T.

UT 11:47 - powering off LBCs as clouds are visible and so we won't do twilgiht flats

UT 12:00 - mask is now back in FPU thanks to Dave T. - error was a post-condition violation. On "release mask" one of the strain gauges stayed ON, which triggered alert.

UT 12:07 - closing up

Summary

  1. We got the initial acquisition completed for AGW3 and PEPSIPOL, and are 90% finished with setting the rotation angle. The next step is to try off-axis acquisition and guiding.
  2. Only part of the AGW Transform was done
  3. N375 G210 test exposures taken
  4. New Pointing model from Dave T. used - worked relatively well

-- %USERSIG{BarryRothberg - 2017-09-08}%
Topic revision: r6 - 09 Sep 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback